### v1

# LLMShot: Reducing snapshot testing maintenance via LLMs

Abstract
Snapshot testing has emerged as a critical technique for UI validation in modern software development, yet it suffers from substantial maintenance overhead due to frequent UI changes causing test failures that require manual inspection to distinguish between genuine regressions and intentional design changes. This manual triage process becomes increasingly burdensome as applications evolve, creating a need for automated analysis solutions. This paper introduces LLMShot, a novel framework that leverages vision-based Large Language Models to automatically analyze snapshot test failures through hierarchical classification of UI changes. To evaluate LLMShot’s effectiveness, we developed a comprehensive dataset using a feature-rich iOS application with configurable feature flags, creating realistic scenarios that produce authentic snapshot differences representative of real development workflows. Our evaluation using Gemma3 models demonstrates strong classification performance, with the 12B variant achieving over 84% recall in identifying failure root causes while the 4B model offers practical deployment advantages with acceptable performance for continuous integration environments. However, our exploration of selective ignore mechanisms revealed significant limitations in current prompting-based approaches for controllable visual reasoning. LLMShot represents the first automated approach to semantic snapshot test analysis, offering developers structured insights that can substantially reduce manual triage effort and advance toward more intelligent UI testing paradigms.
Index Terms:
snapshot testing, software testing, large language models, user interface testing, automated test analysis
I Introduction
Modern software development cycles demand rapid iteration and frequent code changes, creating significant challenges for development teams who must balance software quality while keeping development cycles short [1]. Over the years, many tools and techniques have been developed to help engineers test software systematically at various levels.
Snapshot testing offers a practical approach for user interface (UI) components by capturing the current output of an application and comparing it against future runs to detect changes [2, 3]. Unlike unit and functional tests that verify correctness, snapshot tests simply check whether anything has changed from a previous known state, making them particularly useful for catching unintended modifications during development. Currently used in several large software companies [4], snapshot testing is adopted due to its ease of implementation, allowing developers to write more tests and consequently prevent regressions.
Despite its benefits, snapshot testing faces significant maintenance challenges. Like other GUI testing techniques [5, 6, 7], current approaches frequently fail to comprehend semantic information [8, 9]. Snapshot tests focus solely on identifying visual differences without providing actionable insights, instead only indicating that changes have occurred without meaningful context. When failures occur, as illustrated in Figure 1, developers must triage each failure to determine whether it represents a regression or reflects intentional changes. Current snapshot testing practices suffer from two key limitations: (1) they do not provide detailed explanations of failures, and (2) they produce frequent false positives from minor, acceptable changes [3].
Figure 1: Two Example Diffs for the PADDING_CHANGE Test Failure Class
While recent research explored smarter visual oracles through tools like DiffDroid [10], Owl Eyes [11], and GLIB [12] that employ deep learning classifiers for visual defects, these approaches fail to contextualize the changes. Existing guidelines recommend detailed descriptions for snapshot tests [3], but current tools provide no automated support for generating such explanations.
To address these challenges, we introduce LLMShot, a framework that leverages vision-based Large Language Models (LLMs) to automate snapshot test failure analysis. LLMShot employs Gemma3 models to categorize, quantify, and explain visual discrepancies using a hierarchical classification scheme. Unlike existing pass/fail approaches, our framework provides semantic understanding of UI changes, distinguishing between genuine regressions and intentional design modifications.
Our evaluation on 17 test cases demonstrates that LLMShot achieves strong classification performance, with the Gemma3:4b model reaching 76.47% hit rate and 66.67% F1-score. However, our exploration of ignore mechanisms revealed significant limitations in current prompting-based approaches.
The main contributions of this work are:
  • LLMShot, an intelligent snapshot analysis framework that performs semantic analysis and systematic categorization of visual discrepancies using a hierarchical classification scheme. This represents the first automated approach to snapshot test failure analysis reported in the academic literature, with a pipeline designed for CI/CD workflow integration.
  • A comprehensive dataset [^1] https://zenodo.org/records/15876540 with hierarchical ground truth labels for UI failure categorization, created using a feature-rich iOS demo app that demonstrates real-world features posing challenges for snapshot tests.
II Background and Related Work
II-A Visual GUI and Snapshot Testing
Early automated GUI testing focused on maximizing coverage through event-sequence generation and model-based exploration. Tools like Sapienz [5] used evolutionary algorithms for code coverage and crash discovery, while Stoat [6] built window-transition graphs to reduce redundant paths and combinatorial reducers [13] minimized event combinations for efficiency. ComboDroid [7] combined real-world traces with automated exploration for cost-effective fault detection. Video2Scenario (V2S) [14] automated test script generation through screen recording analysis using neural object detection, reproducing 89% of user actions. CAPdroid [15] enhanced V2S with improved detection and input validation while generating step descriptions from snapshots. More recently, WebExplor [16] and QExplore [17] formulated navigation as reinforcement learning tasks, improving lines executed, screens reached, and crashes detected beyond simple crawling. However, one notable gap in this body of work is the detection of visual regression that does not affect functional behavior but may still significantly impact user experience.
Lately, snapshot testing has gained popularity for validating UI consistency by comparing rendered states with baselines. However, it remains largely understudied in the literature. Bui and Rocha [18] compiled a JEST dataset from GitHub projects, capturing static attributes like stars, issues, and test files, but dynamic test execution remains unexplored. Cruz et al. [3] performed a grey literature review highlighting snapshot testing’s ease of use and regression-prevention benefits, while also noting its fragility and tendency toward false positives.
To address this, research explored smarter visual oracles. DiffDroid [10] executes Android apps across 130+ device profiles to flag inconsistencies with high precision, while dVermin [19] uses metamorphic approaches for layout issues under accessibility settings. Tools like Owl Eyes [11], Nighthawk [20], and GLIB [12] employ deep learning classifiers for visual glitches, achieving impressive precision (0.85 for mobile apps, 0.99 for game sprites) through synthetic data augmentation. However, while these tools can detect a wide range of visual defects, they fail to contextualize the changes, leaving developers without explanations for why the discrepancy occurred or whether it was intentional, which is essential in snapshot testing.
II-B LLM-based Semantic Analysis for UI Verification
LLMs have introduced unprecedented potential for semantic analysis in software engineering, particularly for bug detection and UI verification. GPTDroid [8] frames testing as dialogue, achieving 32% improved activity coverage and 31% more bugs in 93 Android apps while helping developers fix 35 issues. VisionDroid [21] extends this to multimodal tasks, combining OCR with vision-language models to detect complex functional bugs like discrepancies between totals and itemized amounts.
While these advances represent significant progress, most LLM-based testing tools focus on bug detection and test generation, leaving snapshot testing largely unexplored. Traditional snapshot tests suffer from false positives or intentional UI modifications flagged as failures, accounting for up to 8.2% of commits [2] and requiring significant manual review effort. Existing guidelines recommend writing snapshot tests with "detailed and clear descriptions" and treating snapshots with the same care and discipline as code [3]. However, current tools provide no support for generating these detailed descriptions automatically for snapshot failures. LLMShot aims to alleviate this burden by using LLMs to semanticize the snapshot validation process, making it easier to distinguish between expected changes and true regressions.
III Dataset Creation
A key goal in snapshot testing is developing semantic understanding to distinguish intentional feature updates from genuine regressions. This requires realistic scenarios with meaningful explanatory labels beyond binary pass/fail results. While Bui et al. [18] developed what is, to the best of our knowledge, the first and only snapshot-specific dataset to date derived from 686 GitHub repositories, their work focuses on static analysis without executing tests or analyzing actual snapshot content, lacking infrastructure for test execution.
To address these limitations, we built our dataset using a real application with feature flags corresponding to realistic development scenarios: user privilege changes, content updates, and styling modifications. This generates authentic snapshot differences by toggling actual application features developers commonly modify—scenarios that fail snapshot tests but aren’t genuine regressions. Our dataset provides ground truth labels for change categories (Table I) classifying UI modifications from feature flag activation: visual changes (color, padding), content updates (text, data), structural modifications (layout), and behavioral changes (animations, semantic state).
TABLE I: Snapshot Test Failure Categories and Descriptions
Category Description
COLOR_CHANGE Different color values due to styling updates.
PADDING_CHANGE Change in margins/spacing; layout shifts.
CONTENT_CHANGE Different content (e.g. image) with changed meaning.
LAYOUT_CHANGE Components repositioned/resized, changing structure.
TEXT_CHANGE Text string changed, altering displayed message meaning.
ANIMATION_PHASE Snapshot taken mid-animation, showing intermediate state.
ANIMATION_CHANGE Change in animation duration/timing alters output.
SEMANTIC_CHANGE Behavior changed (toggle on/off) without layout/text change.
UNKNOWN_<T> Change not from above categories; T names new reason.
For each record, we run twice, once with default settings and once with feature flags enabled, capturing snapshots at key UI states. This produces baseline and modified images per record. Each record contains: reference.png (baseline snapshot), failure.png (modified snapshot), diff.png (visual diff), and metadata.json (comprehensive context).
IV Methodology
The primary objective is distinguishing genuine UI regressions from intentional design changes and providing explanatory analysis for failing snapshot tests. Our approach employs LLMs to semantically analyze UI differences and categorize change types.
IV-A Core LLMShot
LLMShot processes snapshot test failures by leveraging vision-enabled LLMs. The core prompt is shown in Listing 1. Given a tuple of reference image, failure image, pixel-level diff, and predefined possible categories from Table I, LLMShot produces structured analysis outputs including change classification, quantitative measurements, affected UI elements, and natural language explanations.
Compare three images:
1. REFERENCE: {reference} --- correct baseline image
2. FAILURE: {failure} --- current test image
3. DIFF: {difference} --- pixel-level diff image
Analyze the failure by examining ALL three images carefully.
POSSIBLE CATEGORIES: {categories}
Select ALL applicable categories. Create new ones as "UNKNOWN-<reason>".
DEFINITIONS:
- Pixel-wise diff (0-1): 0 = identical, 1 = completely different
- Semantic diff (0-1): 0 = all components present/correct, 1 = all missing/wrong
Respond with valid JSON only:
"reasons": ["category1", "UNKNOWN-X"],
"pixel_diff": 0.XX,
"semantic_diff": 0.XX,
"elements": ["specific UI component names"],
"explanation": "Clear, concise analysis of key differences"
Listing 1: Core LLM prompt framework of LLMShot
Change classification assigns label(s) from the possible categories or, if the LLM identifies changes that do not fit predefined categories, it can create new categories by giving a descriptive name for the newly identified change type.
The quantitative measurements include a pixel-wise difference score in a 0-1 scale, where 0 indicates identical images and 1 represents completely different images, and a semantic difference score in a 0-1 scale, where 0 means all expected components are present and correct, and 1 indicates all components are missing or incorrect. The affected UI elements output provides a list of specific component names that exhibit changes, while the natural language explanations offer an analysis of the key differences observed between the reference and failure snapshots.
IV-B Extension: Ignore Category Mechanism
In real-world scenarios, developers often need to ignore certain expected failure categories while focusing on others that indicate actual regressions. Developers can either ignore accepted failures identified by core LLMShot or directly specify ignore categories upfront, with the latter approach potentially enhancing the LLM’s sensitivity to detecting other categories. To address this need, we develop an ignore mechanism as an extension to LLMShot that allows developers to specify acceptable change categories, enabling focused analysis on unexpected differences that may indicate genuine issues.
The implementation modifies the core prompt (Listing 1) to instruct the LLM to disregard specified categories while maintaining sensitivity to other modifications, as shown in Listing 2.
{core_prompt}
IGNORE the following aspect of the differences: {ignore_reason}
This difference is acceptable; focus on other differences that might exist.
Listing 2: Ignore reason template for LLM
V Setup
V-A Dataset Construction
Our implementation uses a SwiftUI-based iOS application with tab-based architecture across Home, Profile, and Settings views, employing MVVM patterns with centralized state management. The app includes UI elements that change based on user state and active feature flags: content visibility, layout variations, and dynamic animations. We use runtime-configurable feature flags where each activates specific UI variations similar to code modifications. The combination of user states (free vs. premium) and feature flags creates diverse scenarios—the same flag produces different visual outcomes depending on user interface context, enabling comprehensive analysis with semantic understanding of UI changes.
We achieve snapshot capture by running tests using XCTest and XCUITest, extending default behaviors to support scrolling, element interactions, and app launching with specific configurations. Tests were executed using Xcode 16.3 on iPhone 15 Pro iOS 17.2. We use the swift-snapshot-testing library [22] to capture screenshots, running each test twice as described in the methodology.
We run a total of 17 tests with 19 different flags enabled. Two tests are multi-flag, having two flags enabled simultaneously. The distribution of flag categories across the 17 tests shows that COLOR_CHANGE is the most prevalent type of UI discrepancy, occurring in 5 tests. Three categories—PADDING_CHANGE, CONTENT_CHANGE, and LAYOUT_CHANGE—each appear in 3 tests. TEXT_CHANGE and ANIMATION_PHASE each occur in 2 tests, while ANIMATION_CHANGE is the least frequent, appearing in only 1 test.
The dataset exhibits a mean pixel difference of 0.053 ± 0.054 (mean ± std) across all test cases, calculated using Equation 1. This indicates that, on average, the reference and failure images differ by approximately 5.3% at the pixel level. The substantial variation across test cases demonstrates dataset diversity. We quantify image-level differences using a normalized pixel-wise difference score:
$D_{p}=\frac{1}{N}\sum_{i=1}^{N}\frac{|R_{i}-F_{i}|}{255}$ (1)
where $N$ is the number of pixels, $R_{i}$ and $F_{i}$ are RGB values at pixel $i$ in reference and failure images, respectively.
V-B LLMShot
We evaluate LLMShot using Gemma3 [23] models in 4B and 12B parameter variants with a temperature of 0.1 through the Ollama framework [24]. The core prompt framework (Listing 1) guides the LLM to analyze reference, failure, and diff images systematically, producing structured JSON responses.
V-C Evaluation Metrics
Performance assessment combines classification and difference estimation metrics.
Classification metrics include Test Case Hit Rate (percentage of cases with at least one correct ground truth category), Recall (proportion of ground truth labels correctly predicted), Precision (proportion of correct model predictions), F1-Score (balanced precision-recall measure penalizing false positives), Average Labels Per Test (model’s prediction tendency), and Unknown Rate (cases defaulting to UNKNOWN_<T> classifications).
For visual discrepancy quantification, Ground Truth Pixel Difference uses deterministically calculated values (Equation 1), Predicted Pixel Difference represents LLM estimates, Error Pixel Difference measures absolute prediction errors calculated as $|D_{p}^{predicted}-D_{p}^{ground\_truth}|$, and Semantic Difference captures model-generated perceptual dissimilarity scores.
For ignore mechanism evaluation, Ignore Compliance Rate measures the percentage of tests where the specified ignore category was not predicted (indicating successful adherence to ignore instructions), and classification metrics are calculated after removing ignored categories from both predictions and ground truth to assess performance on remaining categories.
VI Evaluation
VI-A RQ1: How does LLMShot perform in Failure Classification?
The primary objective is to establish whether visual language models can accurately categorize different types of UI changes in snapshot test failures, providing developers with structured explanations of detected differences. To evaluate LLMs’ automated snapshot failure analysis capability, we applied the core LLMShot prompt framework to our dataset.
Table II shows LLMShot’s classification performance across model sizes. The larger gemma3:12b model outperformed the smaller gemma3:4b model across all metrics, achieving over 80% Hit Rate and Recall. However, Precision was notably lower for both models, indicating some incorrect predictions. F1-Scores of 74.42% and 66.67% respectively demonstrate balanced performance when accounting for false positives. Both models showed conservative prediction behavior, averaging fewer than 1.5 labels per test.
Finding 1: Our results indicate that LLMShot demonstrates strong potential in automatically classifying snapshot test failure root causes on our dataset.
TABLE II: Performance of Core LLMShot
Model gemma3:4b gemma3:12b
Hit Rate (%) 76.47 82.35
Recall (%) 78.95 84.21
Precision (%) 57.69 66.67
F1-Score (%) 66.67 74.42
Avg. # label/test 1.53 $\pm$ 0.50 1.41 $\pm$ 0.49
Unknown Rate (%) 11.76 5.88
Predicted 0.126 $\pm$ 0.186 0.066 $\pm$ 0.058
Error 0.095 $\pm$ 0.161 0.059 $\pm$ 0.048
Semantic Diff. 0.921 $\pm$ 0.046 0.079 $\pm$ 0.079
  • Avg. # label/test: Average Number of Predicted Labels per Test
  • Hit Rate: Test Case Hit Rate
  • Semantic Diff.: Semantic Difference
VI-B RQ2: How does the ignore-category mechanism affect performance of LLMShot?
In real-world development scenarios, developers must selectively ignore failure categories corresponding to intentional changes while maintaining detection of categories that indicate actual regressions. One approach is to use core LLMShot, ignore the intended classes, and highlight other classes from its results, since LLMs can provide multiple labels. The other approach is to use an extension of the ignore category mechanism. Here, we investigate whether LLMShot can selectively ignore specified UI change categories while maintaining sensitivity to other change types.
To evaluate this extension on our dataset, we follow two strategies: "Ignore From Analysis" runs the core LLMShot once and then runs the extension by instructing the LLM to ignore the first category identified by core LLMShot, and "Ignore From Ground Truth" runs the extension while ignoring the first ground truth category. In both approaches, if LLMShot outputs the ignored category, it is considered a false positive. When the ignored category represents the only correct classification, the expected LLMShot response is to return an empty set indicating no other categories found, and this is considered a hit.
For this experiment, we used the smaller Gemma3:4b model due to its suitability for CI/CD deployment and practical implementation constraints. Table III reveals significant performance degradation across all ignore strategies compared to core LLMShot. "Ignore From Analysis" performed worst, with Hit Rate dropping to 18.75%, Recall to 28.57%, and F1-Score plummeting to 13.79%. "Ignore From Ground Truth" showed better recall but maintained the same low precision (9.09%), resulting in only marginal F1-Score improvement to 16.67%. Both ignore strategies demonstrated poor compliance rates (31.25% and 35.29% respectively), indicating LLMShot frequently predicted ignored categories despite explicit instructions. Additionally, pixel difference errors increased substantially (0.157±0.244 and 0.287±0.339 vs 0.095±0.161 baseline), and unknown classification rates rose, particularly for "Ignore From Analysis" (25% vs 11.76% baseline).
Finding 2: Current LLM-based ignore mechanisms struggle severely with selective change filtering, showing substantial performance degradation across all classification metrics and poor instruction compliance compared to core LLMShot analysis.
[…]
Metric Core LLMShot IFA IFGT
Hit Rate (%) 76.47 18.75 41.18
Recall (%) 78.95 28.57 100.00
Precision (%) 57.69 9.09 9.09
F1-Score (%) 66.67 13.79 16.67
Avg. # label/test 1.53$\pm$0.50 1.38$\pm$0.60 1.29$\pm$0.96
Unknown Rate (%) 11.76 25.00 11.76
IC Rate (%) 31.25 35.29
Predicted 0.126$\pm$0.186 0.194$\pm$0.255 0.332$\pm$0.326
Error 0.095$\pm$0.161 0.157$\pm$0.244 0.287$\pm$0.339
Semantic Diff. 0.921$\pm$0.046 0.894$\pm$0.050 0.891$\pm$0.060
  • Avg. # label/test: Average Number of Predicted Labels per Test
  • IC: Ignore Compliance
  • IFA: Ignore from Analysis
  • IFGT: Ignore from Ground Truth
VII Discussion
Our evaluation demonstrates that LLMShot successfully addresses a fundamental gap in snapshot testing by providing semantic understanding of UI changes. Unlike existing visual testing tools such as DiffDroid [10] and Owl Eyes [11] that focus primarily on detection, LLMShot offers explanatory analysis that helps developers distinguish between intentional design changes and genuine regressions. The strong classification performance achieved by both Gemma3 variants indicates that vision-language models possess sufficient visual reasoning capabilities for practical snapshot test analysis.
However, our exploration of ignore mechanisms reveals fundamental limitations in current prompting-based approaches for controllable visual reasoning. The substantial performance degradation when attempting selective change filtering suggests that simply instructing LLMs to ignore specific categories is insufficient. This finding highlights a broader challenge in LLM-based systems: while these models excel at comprehensive analysis, achieving fine-grained control over their attention and reasoning processes remains difficult.
Our dataset construction approach, using feature flags to generate realistic UI variations, demonstrates the importance of authentic test scenarios for evaluating visual analysis systems. Unlike synthetic datasets or isolated visual defects, our methodology captures the nuanced differences that occur in real development environments.
Internal Validity: Our evaluation uses a limited number of test cases, which may affect generalizability. However, we ensured dataset diversity through real-world scenarios with varied UI modifications, making our sample representative of practical development contexts.
Construct Validity: Our predefined categories may not encompass all UI change types. We address this through the UNKNOWN label mechanism and focus on common failure scenarios encountered in development practice. Author-labeled ground truth could benefit from expert validation.
Reliability: LLM inherent randomness may produce varying outputs across runs. We mitigate this through low temperature settings and consistent hyperparameters, ensuring reproducible results.
VIII Conclusion
This paper presented LLMShot, the first automated framework for semantic snapshot test analysis that leverages vision-based Large Language Models to classify and explain UI test failures. Our comprehensive evaluation demonstrates that LLMShot can achieve over 84% recall in identifying failure root causes, representing a significant advancement toward intelligent UI testing.
The key contributions include: (1) a novel application of vision-language models to snapshot testing challenges, (2) a realistic evaluation dataset constructed through feature flag variations, and (3) empirical evidence of both the potential and limitations of current LLM-based visual reasoning approaches. While our ignore mechanism exploration revealed constraints in controllable reasoning, the core classification capabilities show promise for reducing manual triage effort in real development workflows.
Future work should investigate how LLM-generated explanations affect developer cognitive load and decision-making efficiency in real-world development scenarios, including whether structured explanations reduce triage time compared to traditional manual inspection. Additionally, exploring the quality and interpretability of LLM explanations through developer studies could reveal how explanation verbosity and technical detail impact practical utility in production environments. LLMShot establishes a foundation for advancing snapshot testing from binary pass/fail systems toward intelligent, explanatory analysis that supports informed developer decision-making.